OpenStudy (anonymous):

A 2-lb bag of brown rice costs six dollars. How much would you expect to pay for a 5-lb bag of brown rice, assuming there is no discount for buying a larger bag?

OpenStudy (anonymous):

Turn the problem into a ratio. So you would have $6/2lb=x/5lb. You then either use algebra or the cross multiplication and divide method to solve. Cross multiply and divide method: You multiply $6 times 5lb to get 30. Then you divide by 2lb to get $15 as your answer. To solve with algebra: Multiply both sides of the equation by 5lb. (5lb)($6)/2lb=(x)(5lb)/5lb This reduces to 30/2=x So x= $15